We are given the demand function P =100-5x at Price P=20 Find of elasticity demand?


1
Expert's answer
2021-03-01T10:18:53-0500

"P=100-5x"

"20=100-5x"

"5x=80"

"x=16"

again, "5x=-P+100"

"x=\frac{-P}{5}+\frac{100}{5}"

"E_d=\frac{dx}{dP}\times \frac{P}{x}=\frac{-1}{5}\times \frac{20}{16}=-0.25"